OpenStudy (rock_mit182):

On earth, atmospheric pressure at sea level is 101.3 kPascals. at Bogota 74.6kP respectively. nitrogen exerts 79% of the atmospheric pressure, to calculate earth’s atmospheric pressure, EP, relating pluton’s atmospheric pressure, PP, you could use the relation;

OpenStudy (rock_mit182):

\[a) EP = \frac{ PP }{ 10130 }\] \[b) EP = \frac{ 10130 }{ PP }\] \[C) EP = PP*10130\] \[D)EP = PP +\frac{ 10130 }{ PP }\]

OpenStudy (rock_mit182):

@abb0t @aaronq @Frostbite @JFraser.

OpenStudy (aaronq):

pluton's?

OpenStudy (aaronq):

what does that have to do with earth's atm pressure?

OpenStudy (rock_mit182):

"pluto's " dwarft planet

OpenStudy (rock_mit182):

Idk, is just the way the problem was made

OpenStudy (aaronq):

yeah, i don't know how that makes any sense. sorry

OpenStudy (rock_mit182):

@halorazer could you help me out

OpenStudy (anonymous):

Is that the entire question? If so, english is your teacher's second language.

OpenStudy (rock_mit182):

lol it is

OpenStudy (rock_mit182):

Do you know how to solve this ?

OpenStudy (anonymous):

I probably could if it was worded better. I have a feeling it has something to do with proportions though.